OpenStudy (anonymous):

LCM x^2+9x, x^2-9x whats the LCM???

OpenStudy (anonymous):

First factor each. Can you factor x^2+9x ? Can you factor x^2-9x ?

OpenStudy (anonymous):

no so would it be zero

OpenStudy (anonymous):

oh okay i got it now thanks

OpenStudy (anonymous):

Im still stuck would it be x^2(9-9)

OpenStudy (anonymous):

x^2+9x = x(x + 9) x^2-9x = x(x - 9) So each expression has 2 factors. Do they share any "common factors"?

OpenStudy (anonymous):

just x

OpenStudy (anonymous):

yikes. nevermind! that's the HCF. The LCM is x(x + 9)(x - 9)

OpenStudy (anonymous):

thank you alot

OpenStudy (anonymous):

sure. no problem.